A fixed mass of gas occupies 1000cm3 at 0•C If it is heated at constant pressure to 100•C, calculate its new volume


1
Expert's answer
2021-03-31T09:38:42-0400

Answer

For constant pressure

"\frac{V_2}{V_1}=\frac{T_2}{T_1}\\"


"V_2=1000\times \frac{373}{273}"


"V_2=1366.3cm^3"